Trent can deliver his newspapers in 60 minutes. Lois delivers her newpapers in 40 minutes on the same route. How long would it take them to deliver the newspapers if they work together? ________minutes?
Answer by stanbon(75887)

Trent can deliver his newspapers in 60 minutes. Lois delivers her newpapers in 40 minutes on the same route. How long would it take them to deliver the newspapers if they work together? ________minutes?
----------------------
Trent DATA:
Time = 60 min/job ; rate = 1/60 job/min
---------------------
Lois DATA:
Time = 40 min/job ; rate = 1/40 job/min
-------------------------
Together DATA:
Time = x min/job ; rate = 1/x job/min
--------------------------
EQUATION:
rate + rate = together rate
1/60 + 1/40 = 1/x
Simplify the left side as follows:
100/2400 = 1/x
x = 2400/100
x = 24 minutes (time to deliver the papers if they work together)
===================
Cheers,
Stan H.